This Quiz: 14 pts poss Question Help ng the re ulity a nha. gthat is reported t increase the umber of hours a sleep Construct a 90% o tio nce nerval Totest the ene benessa anew n patients get during the night, researchers randomly select 13 patents and record the number of hours of sleep each gets with and without the new drug Patient Sleep without the Sleep using the drug 6615664 53 71 40 58 33 5461 56 50 34 The condence interval is□<h□ Type an integer or a decimal Round to two decimal places as needed)
Answer #1

From the given data,values of d are got as follows:

- 22, - 1.2, - 2.5, - 1.7, - 2.6, - 1.8, - 2.2, - 1.2, - 1.9, - 2.4, - 2.3, - 2.3, -0.9

From the d values, the following statistics are calculated:

.n =13

ar{d}- 25.2/13 = - 1.9385

sd = 0.5485

SE = sd/sqrt{n}

= 0.5485/ИЗ = 0.1521

alpha = 0.10

ndf = 13 - 1 =12

From Table, critical values of t = pm 1.7823

So,

Confidence Interval:

- 1.9385 pm (1.7823 X 0.1521)

= - 1.9385 pm 0.2711

= (- 2.2096, - 1.6674)

So,

the confidence interval is:

- 2.21 < mu _{d} < - 1.67